# A Reproducible Framework for Neural Topic Modeling in Focus Group Analysis

### v1

@@ Abstract
Focus group discussions generate rich qualitative data but their analysis traditionally relies on labor-intensive manual coding that limits scalability and reproducibility. We present a rigorous, reproducible computational framework for applying neural topic modeling to focus group transcripts, addressing fundamental methodological challenges: hyperparameter sensitivity, model stability, and validation of interpretability. Using BERTopic applied to ten focus groups exploring HPV vaccine perceptions in Tunisia (1,076 utterances), we conducted systematic evaluation across 27 hyperparameter configurations, assessed stability through bootstrap resampling with 30 replicates per configuration, and validated interpretability through formal human evaluation by three domain experts. Our analysis demonstrates substantial sensitivity to hyperparameter choices and reveals that metric selection for stability assessment must align with analytical goals. A hierarchical merging strategy (extracting fine-grained topics for stability then consolidating for interpretability) effectively navigates the stability-coherence tradeoff, achieving coherence of 0.558 compared to 0.539 for direct extraction. Human validation confirmed topic quality with very good inter-rater reliability (ICC = 0.79, weighted Cohen’s kappa = 0.578). Our framework provides practical guidelines that researchers can adapt to their own qualitative research contexts. All code, data processing scripts, and evaluation protocols are publicly available to support reproduction and extension of this work.
@@ 1 Introduction
Topic modeling has emerged as a powerful computational approach for discovering latent thematic structures in large text corpora. Originally developed for document organization and information retrieval, topic modeling techniques have found increasing application in public health research, where understanding discourse patterns can inform intervention strategies and policy decisions. The COVID-19 pandemic accelerated this trend, with numerous studies applying topic modeling to analyze public sentiment and concerns expressed on social media platforms [qorib_covid-19_2023, figueiredo_topic_2024, yin_sentiment_2022, lyu_covid-19_2021, boon-itt_public_2020]. Similarly, topic modeling has been deployed to characterize discussions about Human Papillomaviruss (HPV) vaccination [du_use_2020, surian_characterizing_2016, rai_detecting_2024], cancer care experiences [steele2025], and mental health support communities [carron-arthur_whats_2016]. These applications demonstrate the growing recognition of topic modeling as a valuable tool for public health surveillance and understanding health-related discourse at scale.
However, the majority of these applications focus on readily available online data sources such as Twitter, Reddit, and online forums. While social media provides unprecedented access to large-scale public discourse, it represents only one facet of health communication. Focus group discussions, a cornerstone methodology in qualitative health research, offer structured settings where participants engage in moderated conversations that elicit deeper, contextualized perspectives on health beliefs, barriers, and decision-making processes. Despite their methodological importance, focus group transcripts have received limited attention in the topic modeling literature. This gap likely reflects both the proprietary nature of focus group data and the traditional emphasis on human-coded qualitative analysis. Yet focus groups generate rich textual data that could benefit from computational approaches, particularly when research objectives include systematic synthesis across multiple sessions or geographic contexts.
The application of topic modeling to qualitative research raises important methodological challenges. Focus group transcripts are typically small, heterogeneous corpora with non-uniform conversational structure, which can push topic models outside of the conditions under which they are commonly benchmarked. In such low-data regimes, model outcomes can be highly sensitive to design choices. At the same time, topic modeling algorithms are inherently stochastic, which means that seemingly minor configuration changes can lead to very different topic structures which raises questions about the reproducibility and stability of results, especially when findings inform public health interventions. Recent methodological reviews have highlighted widespread issues with reproducibility in topic modeling research [hagg2022] and emphasized the multidimensional nature of model quality [hosseiny2023].
@@
This paper presents a rigorous, reproducible framework for applying BERTopic [grootendorst2022], a neural topic model, to focus group discussion transcripts. Using data from ten focus groups exploring HPV vaccine perceptions in Tunisia, we address three methodological objectives: (1) systematic exploration of hyperparameter configurations to understand their influence on model outcomes, (2) assessment of topic stability through bootstrap resampling to quantify robustness, and (3) validation of topic interpretability through human evaluation. Our work makes three primary contributions. First, we demonstrate the feasibility and value of applying transformer-based topic modeling to focus group transcripts. Second, we provide a transparent, reproducible framework for hyperparameter selection that balances multiple quality criteria including stability, coherence, and interpretability. Third, through detailed methodological reporting and comparison with traditional approaches, we contribute to the growing literature on rigorous practices in topic modeling for qualitative health research.
@@
To ensure full reproducibility, we provide complete code, evaluation scripts, and documentation of all modeling decisions.
@@ 2.2 Stability and Quality Assessment
Stability is typically quantified using partition similarity measures such as Normalized Mutual Information (NMI) or Adjusted Rand Index (ARI) [hubert1985, strehl2002]. However, these metrics measure fundamentally different aspects of clustering agreement. ARI quantifies exact pairwise assignment agreement and is sensitive to cluster boundary variations [hubert1985], while NMI measures information structure preservation based on entropy [strehl2002]. Recent systematic evaluations have demonstrated that NMI and ARI can produce different algorithm rankings depending on corpus characteristics [rudiger_topic_2022], indicating that neither metric is universally superior and selection should align with analytical goals. Meaney et al. [meaney2023] provide a comprehensive review of topic model quality indices and demonstrate through empirical application that different evaluation metrics often favor models of different complexity. Hosseiny and Baumer [hosseiny2023] emphasize that stability is multidimensional and that researchers must navigate tradeoffs between stability and other quality criteria.
@@
Semantic coherence, typically measured using the $C_{v}$ metric [roder_exploring_2015], assesses whether topic words form interpretable semantic groups. However, coherence and stability represent distinct and sometimes conflicting quality dimensions [hosseiny2023, koltcov2024]. Koltcov et al. [koltcov2024] demonstrated empirically that models can exhibit high coherence while maintaining poor stability, underscoring the importance of evaluating both dimensions. Human validation remains essential for confirming that computationally optimal topics align with domain-relevant interpretations [nanda2023, steele2025].
2.3 Reproducibility and Applications to Qualitative Research
@@
Reproducibility has emerged as a critical concern in topic modeling research. Hagg et al. [hagg2022] conducted a scoping review of 68 LDA studies in psychological science and found that most provided insufficient methodological detail to reproduce analyses, and that results are highly sensitive to analytic decisions. This highlights the need for transparent reporting of all modeling choices, systematic hyperparameter exploration, and rigorous validation.
Applications to health-related qualitative data remain relatively nascent. Studies have applied topic modeling to patient experience surveys [steele2025], psychotherapy transcripts [lalk2024], and dementia caregiver interviews [he2024], demonstrating its efficiency in identifying thematic patterns while highlighting challenges in domain-informed preprocessing and hierarchical topic organization [dou_hierarchicaltopics_2013]. Several studies have proposed augmented approaches combining computational efficiency with qualitative depth [guetterman_augmenting_2018, chang_accelerating_2021, leeson_natural_2019, crowston_using_2012].
@@
Despite these advances, we could find no prior work explicitly applying topic modeling to focus group discussion transcripts. This gap is notable given the prevalence of focus groups in health research and the potential for topic modeling to facilitate synthesis across multiple sessions. The present study addresses this gap by developing and validating a rigorous BERTopic framework specifically designed for focus group analysis, with transparent reporting of all methodological decisions to support reproducibility and adaptation by other researchers.
@@ 3 Methodology
We apply BERTopic to HPV vaccine perception focus groups conducted in Tunisia. Our pipeline addresses three objectives: (1) systematic hyperparameter exploration, (2) stability assessment through bootstrap resampling, and (3) validation of topic interpretability. To contextualize BERTopic’s performance, we also implement a tuned Latent Dirichlet Allocation baseline for comparison (detailed in Section 3.5). Figure 1 illustrates our methodology pipeline.
Figure 1: Systematic pipeline for topic modeling of focus group transcripts. The workflow begins with transcription and translation, followed by embedding generation. A grid search explores the chosen hyperparameter configurations, evaluating each through bootstrap stability analysis (NMI, ARI) and performance metrics (coherence, topic count, outlier fraction). Multicriteria selection balances stability and interpretability, with final validation by domain experts yielding the interpretable topic structure.
@@ 3.1 Data collection and description
The resulting corpus contained 1,076 speaker turns or utterances, with a mean length of 32.7 words (SD = 25.0). Table 1 presents a sample of annotated and translated speaker turns.
@@
En vérité, les centres de santé de base sont toujours surpeuplés et il y a des dizaines de patients quotidiennement qui … In fact, basic health centers are always overcrowded, with dozens of patients coming in every day… par97339 Medenine 45
Pardon, mais d’après mon expérience personnelle et ce que j’entends, la majorité des gens ont eu des troubles de mémoire… Excuse me, but from my personal experience and what I hear, most people have had memory problems… par39944 Bardo 47
Lorsqu’une femme souffre d’un cancer du col de l’utérus, elle préfère ne le dire à personne. Elle reste silencieuse, car… When a woman suffers from cervical cancer, she prefers not to tell anyone. She remains silent because… par73108 Sousse 43
Pourquoi ne vaccinons-nous que les filles et pas les garçons …. Why do we vaccinate only girls and not boys?… par83355 Jendouba 38
Nous sommes pour l’obligation de vaccination…. We are in favor of mandatory vaccination… par03182 Sousse 22
Je me suis disputée avec mon mari à cause du vaccin. Il travaille à l’hôpital et a essayé de me convaincre quotidienneme… I argued with my husband about the vaccine. He works at the hospital and tried to convince me every day… par49154 Medenine
Table 1: Sample of annotated and translated speaker turns
@@ 3.2 Embedding Generation and BERTopic Implementation Pipeline
We used the sentence-transformers/paraphrase-multilingual-mpnet-base-v2 model[reimers2019] to generate 768-dimensional sentence embeddings for all preprocessed sentences. This model is multilingual and optimized for sentences semantic similarity tasks, making it relevant for topic modeling in French.
Importantly, the embedding was stored and used across all subsequent model configurations, ensuring that any observed differences in topic quality or stability result from hyperparameter variations rather than inconsistencies in semantic representation. This also provides computational efficiency by eliminating redundant embedding calculations during hyperparameter exploration.
@@ 3.3 Hyperparameter Exploration and Stability Analysis
For each of the 27 combinations of parameters that result, we performed the following evaluation protocol:
Bootstrap Stability Analysis:
@@
We first generated 30 bootstrap replicas by sampling 85% of the original sentences with replacement. This sampling fraction provides sufficient data for reliable topic discovery while introducing meaningful variation to assess model robustness. Each bootstrap replicate used a deterministic random seed (base seed + replica number) to ensure reproducible sampling. For each bootstrap replicate, we instantiated fresh UMAP and HDBSCAN models. This is important because fitted dimensionality reduction and clustering models retain internal state that could bias results if reused across different data samples. Each bootstrap model was fitted independently on its respective data sample. We assessed model stability using Normalized Mutual Information (NMI) and the Adjusted Rand Index (ARI), computed between all pairs of bootstrap replicates. For each metric, pairwise scores were computed only for documents appearing in both replicates, and the median value across all comparisons was retained as the stability estimate for each parameter configuration.
Full Dataset Evaluation:
@@
We then fitted a BERTopic model on the complete dataset using the specified hyperparameters. For each fitted model, we computed topic coherence using $C_{v}$ coherence scores, recorded the number of discovered topics, and calculated the proportion of documents assigned to the outlier category.
For each parameter configuration, we recorded the hyperparameter values, bootstrap stability metrics (median NMI and median ARI), and full dataset metrics (coherence score, number of topics, outlier fraction).
3.4 Model Selection and Validation
We evaluated all parameter configurations across the above listed dimensions: stability (measured by bootstrap NMI and ARI), topic quality (coherence scores), coverage (outlier proportions), and granularity (number of topics). Rather than applying rigid filtering criteria, we examined the full parameter space to understand tradeoffs between competing objectives such as stability versus coherence, and granularity versus interpretability.
Our selection approach prioritized configurations that demonstrate strong performance across multiple criteria while producing analytically useful topic granularity. When multiple configurations showed comparable performance, we conducted detailed comparisons between them and final selection was based on coherence performance and coverage after accounting for the intended analytical requirements.
Selected configurations were validated through multiple complementary approaches. For models producing fine-grained topic structures, we performed hierarchical clustering on topic representations using cosine distance and Ward linkage. We examined the dendrogram and merged topics by selecting a cut height that produces a manageable and interpretable number of super-topics. We computed coherence scores for both original and merged topic structures to quantify the impact of hierarchical consolidation.

### v2

@@ Abstract
Focus group discussions generate rich qualitative data but their analysis traditionally relies on labor-intensive manual coding that limits scalability and reproducibility. We present a systematic framework for applying BERTopic to focus group transcripts using data from ten focus groups exploring HPV vaccine perceptions in Tunisia (1,075 utterances). We conducted comprehensive hyperparameter exploration across 27 configurations, evaluating each through bootstrap stability analysis, performance metrics, and comparison with LDA baseline. Bootstrap analysis revealed that stability metrics (NMI and ARI) exhibited strong disagreement (r = -0.691) and showed divergent relationships with coherence, demonstrating that stability is multifaceted rather than monolithic. Our multi-criteria selection framework yielded a 7-topic model achieving 18% higher coherence than optimized LDA (0.573 vs. 0.486) with interpretable topics validated through independent human evaluation (ICC = 0.700, weighted Cohen’s $\kappa$ = 0.678). These findings demonstrate that transformer-based topic modeling can extract interpretable themes from small focus group transcript corpora when systematically configured and validated, while revealing that quality metrics capture distinct, sometimes conflicting constructs requiring multi-criteria evaluation. We provide complete documentation and code to support reproducibility.
@@ 1 Introduction
Topic modeling has emerged as a powerful computational approach for discovering latent thematic structures in large text corpora. Originally developed for document organization and information retrieval [blei2003], topic modeling techniques have found widespread application across diverse text types. Early work focused on structured document collections such as news articles, scientific papers, and web pages [griffiths2004, newman2006]. More recently, researchers have applied topic modeling extensively to social media platforms including Twitter [qorib_covid-19_2023, lyu_covid-19_2021], Reddit [fong2025, yang2025], and online forums [sik2023, davila2023], as well as to customer reviews [yazici2024, liu2023] and open-ended survey responses [guetterman_augmenting_2018]. Recent advances in transformer-based architectures have enabled neural topic models that leverage contextual embeddings rather than traditional bag-of-words representations, offering improved performance for short, noisy text common in social media [grootendorst2022, egger2022].
Despite this breadth of applications, most implementations target large-scale corpora where statistical inference benefits from thousands or tens of thousands of documents. Moreover, standard benchmarks and evaluation protocols emphasize either well-structured documents (e.g., 20 Newsgroups, scientific abstracts) or homogeneous short texts (e.g., tweets about specific events). This leaves open questions about topic modeling performance on small, heterogeneous conversational corpora where documents vary substantially in length and thematic focus, and where corpus size may comprise only hundreds of utterances.
Focus group discussions represent a particularly important yet computationally underexplored text type. Unlike social media posts, which are spontaneous and often brief, focus groups are structured conversations where participants engage in moderated dialogue following a semi-structured protocol. Unlike customer reviews or survey responses, which target specific products or questions, focus groups explore open-ended topics through emergent conversation. Focus groups are widely used across health research, market research, policy analysis, and user experience studies, generating rich textual data that captures nuanced perspectives through natural conversation. Despite their methodological prevalence, focus group transcripts have received limited attention in the computational topic modeling literature, likely reflecting both the proprietary nature of such data and the traditional emphasis on manual thematic coding. Yet when research objectives include systematic synthesis across multiple sessions, geographic contexts, or longitudinal studies, computational approaches could augment traditional analysis by efficiently identifying thematic patterns while preserving opportunities for human interpretation.
@@
This paper presents a systematic, reproducible framework for applying BERTopic [grootendorst2022], a neural topic model, to focus group discussion transcripts. Using data from ten focus groups exploring HPV vaccine perceptions in Tunisia, we address four methodological objectives: (1) systematic exploration of hyperparameter configurations to understand their influence on model outcomes, (2) assessment of topic stability through bootstrap resampling to quantify robustness, (3) comparison with Latent Dirichlet Allocation to contextualize BERTopic’s performance against traditional approaches, and (4) validation of topic interpretability through independent human evaluation by domain experts. While our application focuses on health discourse, the methodological framework and findings have broader implications for topic modeling of small, heterogeneous conversational corpora common across qualitative research domains.
Our work makes three primary contributions to computational text analysis methodology. First, we develop and validate the first systematic framework specifically designed for topic modeling of focus group transcripts, demonstrating that transformer-based approaches can extract interpretable themes from conversational qualitative data when properly configured and validated. Second, we show that BERTopic substantially outperforms LDA for this data type, achieving 18% higher coherence while discovering more interpretable topics, with human validation confirming that computational coherence metrics successfully predict expert judgments. Third, through transparent documentation of all modeling decisions, systematic evaluation across multiple quality dimensions, and rigorous comparison with traditional approaches, we contribute to the growing literature on reproducible practices in qualitative research.
@@
To ensure full reproducibility, we provide complete code, evaluation scripts, hyperparameter configurations, and detailed documentation of all modeling decisions. Code and anonymized data are available at ….
@@ 2.2 Stability and Quality Assessment
Assessing topic model quality requires evaluating multiple dimensions, as different metrics capture distinct aspects of model performance. Greene et al. [greene2014] advocated for bootstrap stability analysis as a criterion for topic model selection, proposing that models producing stable partitions across resampled data represent more reliable thematic structures. Stability is typically quantified using partition similarity measures such as Normalized Mutual Information (NMI) or Adjusted Rand Index (ARI) [hubert1985, strehl2002]. However, these metrics measure fundamentally different aspects of clustering agreement. ARI quantifies exact pairwise assignment agreement and is sensitive to cluster boundary variations [hubert1985], while NMI measures information structure preservation based on entropy [strehl2002]. Recent systematic evaluations have demonstrated that NMI and ARI can produce different algorithm rankings depending on corpus characteristics [rudiger_topic_2022], indicating that neither metric is universally superior and selection should align with analytical goals.
@@
Semantic coherence, typically measured using the $C_{v}$ metric [roder_exploring_2015], assesses whether topic words form interpretable semantic groups. However, coherence and stability represent distinct and sometimes conflicting quality dimensions [hosseiny2023, koltcov2024]. Koltcov et al. [koltcov2024] demonstrated empirically that models can exhibit high coherence while maintaining poor stability, underscoring the importance of evaluating both dimensions. Furthermore, algorithmic metrics frequently conflict with human judgments of topic quality [hoyle2021, chang2009], and different coherence measures can disagree on model rankings [aletras2013]. Meaney et al. [meaney2023] demonstrate that different evaluation metrics often favor models of different complexity, indicating that no single metric captures all aspects of topic quality. These findings underscore the importance of multi-faceted evaluation that includes expert validation alongside computational metrics.
2.3 Applications to Qualitative Research and Reproducibility Challenges
Topic modeling has found increasing application in qualitative research across multiple domains. Substantial work has applied these methods to social media health discourse, including COVID-19 discussions on Twitter [qorib_covid-19_2023] and HPV vaccination debates [du_use_2020]. More recently, researchers have analyzed structured qualitative data including patient experience surveys [steele2025], psychotherapy transcripts [lalk2024], interview data [he2024], and open-ended survey responses [guetterman_augmenting_2018, chang_accelerating_2021]. These applications demonstrate topic modeling’s potential for efficiently identifying thematic patterns in qualitative data while highlighting methodological challenges including domain-informed preprocessing, handling conversational dynamics, and hierarchical topic organization [dou_hierarchicaltopics_2013].
@@ 2.3 Applications to Qualitative Research and Reproducibility Challenges
However, reproducibility has emerged as a critical concern. Hagg et al. [hagg2022] conducted a scoping review of 68 LDA studies in psychological science and found that most provided insufficient methodological detail to reproduce analyses, and that results are highly sensitive to analytic decisions. This challenge is particularly acute for small, heterogeneous qualitative corpora where configuration choices can profoundly affect outcomes. The lack of standardized evaluation protocols and transparent reporting practices undermines confidence in topic modeling findings, especially when results inform practical decisions. These issues highlight the need for systematic hyperparameter exploration, multi-dimensional quality assessment, rigorous validation protocols, and transparent documentation of all modeling choices.
@@
Despite growing applications, we could find no prior work explicitly applying topic modeling to focus group discussion transcripts. This gap is notable: focus groups generate multi-party conversations where meaning emerges through interactive dialogue, differing substantively from both social media (spontaneous, brief posts) and individual interviews (single-voice narratives). The present study addresses this gap by developing and validating a systematic BERTopic framework for focus group analysis, with transparent reporting of all methodological decisions to support reproducibility and adaptation by other researchers.
@@ 3 Methodology
We apply BERTopic to HPV vaccine perception focus groups conducted in Tunisia. Our pipeline addresses three objectives: (1) systematic hyperparameter exploration, (2) stability assessment through bootstrap resampling, and (3) validation of topic interpretability. To contextualize BERTopic’s performance, we also implement a tuned LDA baseline for comparison (detailed in Section 3.6). Figure 1 illustrates our methodology pipeline.
Figure 1: Systematic pipeline for topic modeling of focus group transcripts. The workflow begins with transcription and translation, followed by embedding generation. A grid search explores the chosen hyperparameter configurations, evaluating each through bootstrap stability analysis (NMI, ARI) and performance metrics (coherence, topic count, outlier fraction). Multicriteria selection balances performance metrics, with final validation by domain experts giving the interpretable topic structure.
@@ 3.1 Data collection and description
The resulting corpus contained 1,075 speaker turns or utterances, with a mean length of 32.7 words (SD = 25.0). Table 1 presents a sample of annotated and translated speaker turns.
@@
En vérité, les centres de santé de base sont toujours surpeuplés et il y a des dizaines de patients quotidiennement qui … In fact, basic health centers are always overcrowded, with dozens of patients coming in every day… par97339 Region1 45
Pardon, mais d’après mon expérience personnelle et ce que j’entends, la majorité des gens ont eu des troubles de mémoire… Excuse me, but from my personal experience and what I hear, most people have had memory problems… par39944 Region2 47
Lorsqu’une femme souffre d’un cancer du col de l’utérus, elle préfère ne le dire à personne. Elle reste silencieuse, car… When a woman suffers from cervical cancer, she prefers not to tell anyone. She remains silent because… par73108 Region3 43
Pourquoi ne vaccinons-nous que les filles et pas les garçons …. Why do we vaccinate only girls and not boys?… par83355 Region4 38
Nous sommes pour l’obligation de vaccination…. We are in favor of mandatory vaccination… par03182 Region3 22
Je me suis disputée avec mon mari à cause du vaccin. Il travaille à l’hôpital et a essayé de me convaincre quotidienneme… I argued with my husband about the vaccine. He works at the hospital and tried to convince me every day… par49154 Region1
Table 1: Sample of annotated and translated speaker turns (utterances). Participants names and regions have been anonymized and the dataset doesn’t contain any Personally Identifiable Information
@@ 3.2 Embedding Generation and BERTopic Implementation Pipeline
We used the sentence-transformers/paraphrase-multilingual-mpnet-base-v2 model[reimers2019] to generate 768-dimensional sentence embeddings for all preprocessed sentences. This model is multilingual and optimized for sentence semantic similarity tasks, making it relevant for topic modeling in French.
Importantly, the embeddings were computed once and stored for use across all subsequent model configurations, ensuring that any observed differences in topic quality or stability result from hyperparameter variations rather than inconsistencies in semantic representation. This also provides computational efficiency by eliminating redundant embedding calculations during hyperparameter exploration.
@@ 3.3 Hyperparameter Exploration and Stability Analysis
For each of the 27 resulting parameter combinations, we performed the following evaluation protocol:
@@
Bootstrap Stability Analysis: We generated 30 bootstrap replicates by sampling 85% of the original sentences with replacement. This sampling fraction provides sufficient data for reliable topic discovery while introducing meaningful variation to assess model robustness. Each bootstrap replicate used a deterministic random seed (base seed + replicate number) to ensure reproducible sampling. For each bootstrap replicate, we instantiated fresh UMAP and HDBSCAN models. This is essential because fitted dimensionality reduction and clustering models retain internal state that could bias results if reused across different data samples. Each bootstrap model was fitted independently on its respective data sample.
We assessed model stability using Normalized Mutual Information (NMI) and the Adjusted Rand Index (ARI), computed between all pairs of bootstrap replicates. For each metric, pairwise scores were computed only for documents appearing in both replicates, and the median value across all pairwise comparisons was retained as the stability estimate for that parameter configuration.
@@
Full Dataset Evaluation: We fitted a BERTopic model on the complete dataset using each parameter configuration. For each fitted model, we computed topic coherence using $C_{v}$ coherence scores, recorded the number of discovered topics, and calculated the proportion of documents assigned to the outlier category.
3.4 Multi-Criteria Model Selection
Rather than applying rigid optimization criteria or automated selection rules, we adopted a multi-criteria decision framework balancing four dimensions:
  • Coherence: Semantic interpretability of topics, measured by $C_{v}$ scores
  • Stability: Bootstrap consistency, assessed through NMI and ARI
  • Granularity: Number of non-outlier topics, determining thematic resolution
  • Coverage: Proportion of documents assigned to topics versus outliers
This approach acknowledges that optimal configurations depend on research context: exploratory qualitative research may prioritize coherence, while studies requiring reproducible classifications emphasize stability. By transparently presenting performance across all 27 configurations, we enable informed selection rather than imposing a single optimization objective.
3.5 Human Validation